Find the equation of a line with a slope of 3/5 that passes throught the point (-1,3)
Answer by stanbon(75887)

Find the equation of a line with a slope of 3/5 that passes throught the point (-1,3)
------------------------
Form: y = mx + b
You are told m=3/5 and y = 3 when x = -1; solve for "b".
---------
3 = (3/5)*-1 + b
b = 3 3/5
b = 18/5
-----------------
Equation:
y = (3/5)x + (18/5)
=========================
Cheers,
Stan H.
Reply to stanbon@comcast.net